# Advantage Shaping as Surrogate Reward Maximization:

### v1

@@ 5.2 Forward-Engineering: Advantage Shaping from Surrogate Reward
For $N\gg K$, the $\widetilde{\text{GRPO}}_{K}$ policy-gradient update by [Che+25] is equivalent to an RLOO-style policy gradient update to the surrogate per-example reward $\frac{2}{K}\arcsin\!\big(\sqrt{\rho_{K,\theta}(x,a)}\,\big)\,$, from Eq. (22).
To see this, we work as in Sec. 3, but now with the arcsin-transformed objective. We proceed in three steps: (i) By direct differentiation, the population per-example gradient is
@@
To obtain an empirical version of this update, we (ii) approximate the multiplicative term of the gradient by substituting $\rho_{K,\theta}$ and $\rho_{\theta}$ with their empirical counterparts $\widehat{\rho}_{K}$ and $\widehat{\rho}$, and (iii) approximate the expected 0/1 reward gradient term with the RLOO gradient update from Eq. (16) (recall Eq. (1)). The update then becomes:
@@
This connection invites exploration of alternative variance-stabilizing transformations. The arcsin transform is a VST for the binomial distribution, which aligns with the batched sampling scheme of [Che+25, Sec. 2.2 ]. However, the standard Pass@K estimator (Eq. (12)) is combinatorial and does not strictly follow a binomial distribution. Deriving a VST tailored to this estimator’s distribution, then applying the forward-engineering recipe from Claim 2, could yield a more stable optimization algorithm.
@@
For $N\gg 1$, the (vanilla) GRPO policy-gradient update (Eq. (9)) by [Sha+24] is equivalent to an RLOO-style policy gradient update for the surrogate per-example reward $2\arcsin\!\Big(\sqrt{\rho_{\theta}(x,a)}\,\Big)\,$, where $\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r_{0/1}(y,a)$ is the expected 0/1 reward per example.
@@
6 A Regularized Surrogate Rewards Percpective
Both Pass@K routes examined in Secs. 3 and 4 tilt the learning signal toward the hard cases. [Che+25] argue this is aligned with the exploit/explore roles of the two metrics. Maximizing the 0/1 reward emphasizes exploitation: making individual responses more often correct, even on already-easy examples. In contrast, Pass@K encourages exploration: for a given example, only one of $K$ responses must be correct, so once an example is “solved” in the Pass@K sense, further pushing its 0/1 reward is less valuable than allocating gradient signal to unsolves examples where more search could pay off. Ideally, we want a controlled balance: retain sufficient signal on solved examples while injecting additional search signal on unsolved ones. To achieve this, they propose mixing Pass@K-shaped and vanilla advantages.
@@ 6.1 Combined 0/1 and Pass@K Advantage Shaping
Having observed that Pass@K optimization can aggressively favor difficult examples (small $\widehat{\rho}$) while potentially zeroing out useful signal on easier ones (medium/large $\widehat{\rho}$), [Che+25] propose a simple blend between optimizing the $0/1$ reward and optimizing Pass@K. They combine the Pass@K advantage with the vanilla advantage using the empirical 0/1 reward $\widehat{\rho}$ as a mixing weight. In our notation:
@@ 6.2 Reward Regularization Interpretation of skew-R
An interesting observation potentially supporting the practical value of skew-R: Kimi 1.5 [Tea+25] performs priority sampling in their RL experiments, reweighting each example by $1-\widehat{\rho}$ to allow harder examples to appear more frequently. This corresponds to the advantage shaping of skew-R (Eq. (27)) applied at the example level rather than within gradient computation. Thus, our interpretation as regularized reward maximization provides a direct justification for their empirically motivated practical choice.
@@
This validates our surrogate-reward reverse-engineering framework, but it also highlights a subtle interpretive question about what constitutes “regularization.” The interpretation depends on which performance metric we take as primary:
  • Pass@K-centric lens. If Pass@K ($\rho_{K}$) is treated as the true performance metric, then GRPOK should be viewed as directly maximizing a smooth, monotone surrogate of $\rho_{K}=1-(1-\rho)^{K}$. In this view there is no extra penalty term: we are simply doing surrogate maximization of the evaluation metric itself.
@@ 6.3 An Example of Reward-Level Regularization
The analysis of skew-R revealed that a simple gradient scaling, $(1-\widehat{\rho})\cdot\widehat{G}_{\text{GRPO}}$, of vanilla GRPO admits a population interpretation as policy gradient optimization with a regularizer term added on the ”good” baseline surrogate $\arcsin(\sqrt{\rho_{\theta}})$ of the original GRPO.
@@
To illustrate the use of reward regularization, we show how to optimize a surrogate reward that combines the data-fitting term with an entropy regularizer $H(\rho)=-\rho\log(\rho)-(1-\rho)\log(1-\rho)$ of the binary 0/1 reward. Intuitively, similar to the std-regularization of skew-R, $H(\rho)$ favors examples where the model is still “undecided,” redistributing gradient effort toward those rather than repeatedly polishing ones it already gets right. The overall surrogate reward is
@@
Applying the forward-engineering recipe (differentiate the per-example regularized objective above, then replace population quantities by their empirical estimates $\hat{\rho}$ and plug in the usual low-variance GRPO estimator for $\nabla_{\theta}\rho$; see Appendix B.4) yields a tunable advantage-shaping rule that scales the vanilla GRPO gradient as follows:
@@
Note that unlike Skew-R’s objective in Eq. (27), which is maximized at $\rho_{\theta}=1$, the entropy-augmented objective need not be globally maximized at $\rho=1$ for all $\lambda$. For sufficiently small $\lambda\lesssim 1.5$, the maximizer of $J_{\lambda}$ remains at $\rho=1$, so we are still ultimately pushing toward perfect success. For larger regularization strengths, the entropy term is strong enough that the overall maximizer shifts to some $\rho_{\lambda}<1$, providing a mechanism to penalize perfect fitting of the training data, a stronger form of regularization.
@@ 7.1 Biased vs Unbiased Scalings
In contrast, [Mah+25] apply a biased empirical gradient that estimates the $\mathrm{Fail}@{(K-1)}$ factor as $(1-\widehat{\rho})^{K-1}$. The corresponding advantage-score scaling is (up to constant multiplicative factor $K$)
@@
Note that such biased scaling is also present in advantage-shapes resulting from our forward-engineering of surrogate rewards in Sec. 5. Even our reverse-engineering operates at the population level. Thus, here, it would tell us that both GRPOK (which applies unbiased reweightings to vanilla GRPO) and the method of [Mah+25], both implicitly maximize the same surrogate reward (see Sec. C.1). In the practical regime of few samples though the two methods could lead to different behaviors. While unbiasedness is intuitively a good property, we discuss below regimes where the biased variant can be particularly attractive. It is worth noting that for $K=2$ the two variants are identical as they coincide with skew-R, which as we discussed in Sec. 5 is also effectively used by [Tea+25].
@@ 7.2 Normalization Factors
Up to the min-clip operation, the main difference of GRPO from classical RLOO is normalization with respect to the reward’s standard deviation (Sec. 2). The surrogate reward perspective (Cor. 1) provides an alternative interpretation of this scaling: at the level of the training objective, it maximizes a transformation of the 0/1 reward that stabilizes the variance of its empirical estimator. This interpretation is consistent for $\widetilde{\text{GRPO}}_{K}$, which as we discussed in Sec. 4 was originally designed through advantage shaping in a batched setting where Pass@K is evaluated on $N/K$ batches of the observed responses. More broadly, our forward-engineering of reward surrogates $F(\rho)$ uses RLOO empirical gradients as a robust estimate of the population reward gradient $\nabla_{\theta}\rho$ and scales them by $F^{\prime}(\rho)$ that generalizes the specific GRPO normalization via standard deviation.
On the empirical side, it is worth noting that there is ongoing debate about whether such variance-stabilizing normalizations are necessary for good performance. For example, in support of earlier works [Ahm+24], a recent study by [Kha+25] finds that RLOO is competitive to GRPO, with zero-mean advantages being the main driver behind its success [Xio+25]. This raises an interesting question: Could we theoretically characterize how properties of a surrogate $F$ (e.g., variance stabilization) relate to favorable optimization dynamics? Such an analysis might shed light on if, when and why these normalizations matter in practice.
@@ 7.3 Granular Advantage Shaping
One can imagine even finer-grained shaping. For example, [wenlong_THR] investigate token-level advantage shaping through the lens of simple one-step dynamics of vanilla expected reward maximization: token-level control incentivizes the model to preserve probability mass on unobserved alternative responses to avoid over-concentrating on observed ones as a means of balancing exploitation and exploration. The entropy-based modification of advantages by [Che+25a] is effectively another instance of more granular token-level control introduced to favor exploration. See also [Cui+25] for yet another token-level advantage shaping motivated by detailed empirical investigation of how entropy relates to downstream reasoning performance.
It would be interesting to apply the reverse-engineering perspective to interpret these token-level methods as regularized surrogate rewards. The finer granularity naturally suggests finer-grained regularization at the policy or even model-representation level. Though this would be more challenging to formalize, a common surrogate-reward perspective could provide an alternate lens to actual benefits brought by the various modifications.
7.4 Clipping
Our analysis omits the min-clip operation of GRPO [Sha+24], which is itself inherited from PPO [Sch+17]. This is accurate when GRPO operates in full online mode (single update per model generation), as originally done by [Sha+24]. However, if multiple updates are performed per generation, the GRPO gradient expression in Eq. (9) that our analysis uses is only a proxy. That said, all advantage-shaping methods studied here (e.g., GRPOK, skew-R, entropy-augmented) can be empirically applied by simply substituting the respective advantage scores into the original GRPO formulation with the min-clip operation (or follow-up variants, e.g., [Yu+25]).
@@ 8 Conclusion
Advantage shaping and surrogate reward design are two sides of the same coin. This equivalence mirrors classical supervised learning, where modifying the training objective is analogous to directly modifying the gradient. Both design approaches have proven effective, with a strong parallel in supervised classification for imbalanced data, e.g. [Lin+17, Men+20, YZC21, Kin+21]. At a conceptual level, understanding this equivalence is instructive: irrespective of the design choice—whether starting from advantage shapes or surrogate rewards—analyzing the counterpart can provide valuable insight. Looking forward, while direct parallels to supervised learning should be drawn cautiously, the (regularized) surrogate reward lens may offer a bridge between iterative optimization algorithms, like policy gradients, and the population performance metrics they ultimately aim to optimize.
@@ Acknowledgements
This work was partially funded by the NSERC Discovery Grant No. 2021-03677 and the Alliance Grant ALLRP 581098-22.
Appendix A Auxiliary Technical Results

### v2

@@ 5.2 Forward-Engineering: Advantage Shaping from Surrogate Reward
For $N\gg K$, the $\widetilde{\text{GRPO}}_{K}$ policy-gradient update by [passk_bydance] is equivalent to an RLOO-style policy gradient update to the surrogate per-example reward $\frac{2}{K}\arcsin\!\big(\sqrt{\rho_{K,\theta}(x,a)}\,\big)\,$, from Eq. (22).
To see this, we work as in Sec. 3, but now with the arcsin-transformed objective. We proceed in two steps: (i) By direct differentiation, the population per-example gradient is
@@
(ii) To obtain an empirical version of this update, we (a) approximate the multiplicative term of the gradient by substituting $\rho_{K,\theta}$ and $\rho_{\theta}$ with their empirical counterparts $\widehat{\rho}_{K}$ and $\widehat{\rho}$, and (b) approximate the expected 0/1 reward gradient term with the RLOO gradient update from Eq. (16) (recall Eq. (1)). The update then becomes:
@@
This connection invites exploration of alternative variance-stabilizing transformations. The arcsin transform is a VST for the binomial distribution, which aligns with the batched sampling scheme of [passk_bydance, Sec. 2.2 ]. However, the standard Pass@K estimator (Eq. (12)) is combinatorial and does not strictly follow a binomial distribution. Deriving a VST tailored to this estimator’s distribution, then applying the forward-engineering recipe from Claim 2, could yield a more stable optimization algorithm.
@@
For $N\gg 1$, the (vanilla) GRPO policy-gradient update (Eq. (9)) by [GRPO_original] is equivalent to an RLOO-style policy gradient update for the surrogate per-example reward $2\arcsin\!\Big(\sqrt{\rho_{\theta}(x,a)}\,\Big)\,$, where $\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r_{0/1}(y,a)$ is the expected 0/1 reward per example.
@@
This surrogate reward perspective should be distinguished from GRPO’s original formulation. As presented by [GRPO_original], GRPO maximizes a PPO-style objective designed for the multi-epoch off-policy setting, where the expectation is taken over the old policy. Our surrogate reward, in contrast, identifies the true on-policy population objective that the GRPO gradient ascends in the fully online setting. We detail this distinction in Appendix A, showing why the PPO-style objective cannot be directly simplified to yield this on-policy surrogate.
6 A Regularized Surrogate Rewards Perspective
Both Pass@K routes examined in Secs. 3 and 4 tilt the learning signal toward the hard cases. [passk_bydance] argue this is aligned with the exploit/explore roles of the two metrics. Maximizing the 0/1 reward emphasizes exploitation: making individual responses more often correct, even on already-easy examples. In contrast, Pass@K encourages exploration: for a given example, only one of $K$ responses must be correct, so once an example is “solved” in the Pass@K sense, further pushing its 0/1 reward is less valuable than allocating gradient signal to unsolves examples where more search could pay off. Ideally, we want a controlled balance: retain sufficient signal on solved examples while injecting additional search signal on unsolved ones. To achieve this, they propose mixing Pass@K-shaped and vanilla advantages.
@@ 6.1 Combined 0/1 and Pass@K Advantage Shaping
Having observed that Pass@K optimization can aggressively favor difficult examples (small $\widehat{\rho}$) while potentially zeroing out useful signal on easier ones (medium/large $\widehat{\rho}$), [passk_bydance] propose a simple blend between optimizing the $0/1$ reward and optimizing Pass@K. They combine the Pass@K advantage with the vanilla advantage using the empirical 0/1 reward $\widehat{\rho}$ as a mixing weight. In our notation:
@@ 6.2 Reward Regularization Interpretation of Skew-R
An interesting observation potentially supporting the practical value of skew-R: Kimi 1.5 [team2025kimi] performs priority sampling in their RL experiments, reweighting each example by $1-\widehat{\rho}$ to allow harder examples to appear more frequently. This corresponds to the advantage shaping of skew-R (Eq. (27)) applied at the example level rather than within gradient computation. Thus, our interpretation as regularized reward maximization provides a direct justification for their empirically motivated practical choice.
@@
This validates our surrogate-reward reverse-engineering framework, but it also highlights a subtle interpretive question about what constitutes “regularization.” The interpretation depends on which performance metric we take as the target:
  • Pass@K performance target. If Pass@K ($\rho_{K}$) is treated as the true performance metric, then GRPOK should be viewed as directly maximizing a smooth, monotone surrogate of $\rho_{K}$. In this view there is no extra penalty term: we are simply doing surrogate maximization of the evaluation metric itself.
@@ 6.3 An Example of Reward-Level Regularization
The analysis of skew-R revealed that a simple gradient scaling, $(1-\widehat{\rho})\cdot\widehat{G}_{\text{GRPO}}$, of vanilla GRPO admits a population interpretation as policy gradient optimization with a regularizer term added on the “good” baseline surrogate $\arcsin(\sqrt{\rho_{\theta}})$ of the original GRPO.
@@
To illustrate the use of reward regularization, we show how to optimize a surrogate reward that combines the data-fitting term with an entropy regularizer $H(\rho)=-\rho\log(\rho)-(1-\rho)\log(1-\rho)$ of the binary 0/1 reward. Intuitively, similar to the standard-deviation regularization of skew-R, $H(\rho)$ favors examples where the model is still “undecided,” redistributing gradient effort toward those rather than repeatedly polishing ones it already gets right. The overall surrogate reward is
@@
Applying the forward-engineering recipe (differentiate the per-example regularized objective above, then replace population quantities by their empirical estimates $\hat{\rho}$ and plug in the usual low-variance RLOO estimator for $\nabla_{\theta}\rho$; see Appendix B.5) yields a tunable advantage-shaping rule that scales the vanilla GRPO gradient as follows:
@@
Note that unlike Skew-R’s objective in Eq. (27), which is maximized at $\rho=1$, the entropy-augmented objective need not be globally maximized at $\rho=1$ for all $\lambda$. For sufficiently small $\lambda\lesssim 1.5$, its maximizer remains at $\rho=1$, so we are still ultimately pushing toward perfect success. For larger regularization strengths, the entropy term is strong enough that the overall maximizer shifts to some $\rho<1$, providing a mechanism to penalize perfect fitting of the training data, a stronger form of regularization.
@@ 7.1 Biased vs Unbiased Scalings
In contrast, [sadegh] apply a biased empirical gradient that estimates the $\mathrm{Fail}@{(K-1)}$ factor as $(1-\widehat{\rho})^{K-1}$. The corresponding advantage-score scaling is (up to constant multiplicative factor $K$)
@@
Note that such biased scaling is also present in advantage-shapes resulting from our forward-engineering of surrogate rewards in Sec. 5. Even our reverse-engineering operates at the population level. Thus, here, it would tell us that both GRPOK (which applies unbiased reweightings to vanilla GRPO) and the method of [sadegh], both implicitly maximize the same surrogate reward (see Sec. C.1). In the practical regime of few samples though the two methods could lead to different behaviors. While unbiasedness is intuitively a good property, we discuss below regimes where the biased variant can be particularly attractive. It is worth noting that for $K=2$ the two variants are identical as they coincide with skew-R, which as we discussed in Sec. 5 is also effectively used by [team2025kimi].
@@ 7.2 Normalization Factors
Up to the clip operation, the main difference of GRPO from classical RLOO is normalization with respect to the reward’s standard deviation (Sec. 2). The surrogate reward perspective (Cor. 1) provides an alternative interpretation of this scaling: at the level of the training objective, it maximizes a transformation of the 0/1 reward that stabilizes the variance of its empirical estimator. More broadly, our forward-engineering of reward surrogates $F(\rho)$ uses RLOO empirical gradients as a robust estimate of the population reward gradient $\nabla_{\theta}\rho$ and scales them by $F^{\prime}(\rho)$ that generalizes the specific GRPO normalization via standard deviation.
On the empirical side, it is worth noting that there is ongoing debate about whether such normalizations are necessary for good performance. For example, in support of earlier works [ahmadian2024back], a recent study by [khatri2025art] finds that RLOO is competitive to GRPO, with zero-mean advantages being the main driver behind its success [xiong2025minimalist]. This raises an interesting question: Could we theoretically characterize how properties of a surrogate $F$ (e.g., variance stabilization) relate to favorable optimization dynamics? Such an analysis might shed light on if, when and why these normalizations matter in practice.
7.3 Clipping
Our analysis omits the clip operation of GRPO [GRPO_original], which is itself inherited from PPO [PPO]. This is accurate when GRPO operates in full online mode (single update per model generation), as originally done by [GRPO_original]. However, if multiple updates are performed per generation, the GRPO gradient expression in Eq. (9) that our analysis uses is only a proxy. That said, all advantage-shaping methods studied here (e.g., GRPOK, skew-R, entropy-augmented) can be empirically applied in practical, multi-epoch training regimes by simply substituting the respective advantage scores into the original GRPO formulation with the clip operation (or follow-up variants, e.g., [DAPO, GSPO]). See Sec. A for a detailed discussion.
@@ 7.4 Granular Advantage Shaping
One can imagine even finer-grained shaping. For example, [thr] investigates token-level advantage shaping through the lens of simple one-step dynamics of vanilla expected reward maximization: token-level control incentivizes the model to preserve probability mass on unobserved alternative responses to avoid over-concentrating on observed ones as a means of balancing exploitation and exploration. The entropy-based modification of advantages by [cheng2025reasoning] is effectively another instance of more granular token-level control introduced to favor exploration. See also [cui2025entropy] for yet another token-level advantage shaping motivated by detailed empirical investigation of how entropy relates to downstream reasoning performance.
It would be interesting to apply the reverse-engineering perspective to interpret these token-level methods as regularized surrogate rewards, albeit in the more form of finer-grained regularization at the policy or even model-representation level. Though this would be more challenging to formalize, a common surrogate-reward perspective could provide an alternate lens to actual benefits brought by the various modifications.
@@ 8 Conclusion
Advantage shaping and surrogate reward design are two sides of the same coin. This equivalence mirrors classical supervised learning, where modifying the training objective is analogous to directly modifying the gradient. Both design approaches have proven effective, with a strong parallel in supervised classification for imbalanced data, e.g. [lin2017focal, menon_LA_loss, ye2021procrustean, kini2021label]. At a conceptual level, understanding this equivalence is instructive: irrespective of the design choice—whether starting from advantage shapes or surrogate rewards—analyzing the counterpart can provide valuable insight. Looking forward, while direct parallels to supervised learning should be drawn cautiously, the (regularized) surrogate reward lens may offer a bridge between iterative optimization algorithms, like policy gradients, and the population performance metrics they ultimately aim to optimize.
@@ Acknowledgements
This work was funded by the NSERC Discovery Grant No. 2021-03677 and the Alliance Grant ALLRP 581098-22.
Summary of Notations and Algorithms
Notations.
For the reader’s convenience, we list the core notation used throughout the paper.
$(x,a)$
A problem-prompt ($x$) and reference-answer ($a$) pair.
$\pi_{\theta}(\cdot|x)$
The policy (LLM) parameterized by $\theta$.
NN
Number of IID responses generated per example.
$y_{i}$
The $i$-th response, $y_{i}\sim\pi_{\theta}(\cdot|x)$.
$r_{i}$
The binary 0/1 reward $r_{0/1}(y_{i},a)$.
$\rho=\rho_{\theta}$
The (population) expected 0/1 reward $\mathbb{E}_{y\sim\pi_{\theta}}[r_{i}]$.
$\widehat{\rho}$
The empirical 0/1 reward $\frac{1}{N}\sum_{i=1}^{N}r_{i}$.
$N^{+},N^{-}$
Number of correct ($r_{i}=1$) and wrong ($r_{i}=0$) responses.
$\rho_{K}=\rho_{K,\theta}$
The (population) expected Pass@K reward.
$\widehat{\rho}_{K}$
The unbiased empirical estimator of $\rho_{K,\theta}$.
$\nabla_{i}$
The log-probability gradient $\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$.
$\widehat{\nabla}_{+}$
Average empirical gradient over correct responses: $\frac{1}{N^{+}}\sum_{i:r_{i}=1}\nabla_{i}$.
$\widehat{\nabla}_{-}$
Average empirical gradient over wrong responses: $\frac{1}{N^{-}}\sum_{i:r_{i}=0}\nabla_{i}$.
Appendix A GRPO’s PPO-Style vs Surrogate Reward Objectives
In the original paper, [GRPO_original] present GRPO as optimizing a PPO-style objective [PPO]. This section reconciles that formulation with the surrogate reward objective derived in our Corollary 1.
PPO-style objective.
Following common practice [DAPO, DrGRPO, rlzoo, deng2025effect], we omit the KL regularization term and the $1/|y_{i}|$ length normalization (which is found responsible for length-bias) from the original GRPO objective of [GRPO_original], yielding:
$\mathbb{E}_{\{y_{i}\}_{i\in[N]}\stackrel{{\scriptstyle\text{IID}}}{{\sim}}\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{|y_{i}|}\min(\gamma_{i,t}A_{i},\left[{\gamma_{i,t}}\right]_{1-\epsilon}^{1+\epsilon}A_{i})\Bigg]\,$, (30)
where $\left[{\cdot}\right]_{a}^{b}$ is the clip function at levels $a$ and $b$, and $\gamma_{it}=\pi_{\theta}(y_{it}|x,y_{i,<t})/\pi_{\theta_{\text{old}}}(y_{it}|x,y_{i,<t})$ is the policy ratio between the current model $\theta$ (to be updated) and the model $\theta_{\text{old}}$ that generates the responses. For clarity, further omit the clip operation, as done for example in [GRPO_original, App. A.1.6 ]. Then the above objective simplifies to
$\mathbb{E}_{y_{i}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{|y_{i}|}\gamma_{i,t}\cdot A_{i}\Bigg]\,.$
Note that the expectation here is defined with respect to samples from the old policy $\pi_{\theta_{\mathrm{old}}}$, following the TRPO and PPO frameworks [trpo, PPO]. Thus, when taking gradients, $\pi_{\theta_{\text{old}}}$ is held fixed; only the numerator $\pi_{\theta}$ depends on $\theta$.
Taking the gradient with respect to $\theta$ and using the log-derivative trick, the gradient becomes
$\mathbb{E}_{y_{i}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{|y_{i}|}\gamma_{i,t}\cdot A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(y_{it}|x,y_{i,<t}|x)\Bigg]\,.$ (31)
The fully online setting (corresponding to $\mu=1$ iterations in [GRPO_original, Alg. 1]) implies $\pi_{\theta}=\pi_{\theta_{\mathrm{old}}}$. Then, by $\gamma_{i,t}=1$, this simplifies to
$\mathbb{E}_{y_{i}\sim\pi_{\theta(\cdot|x)}}\Big[\frac{1}{N}\sum_{i=1}^{N}A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\Big]\,.$ (32)
The empirical version of this, $\frac{1}{N}\sum_{i=1}^{N}A_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$, is precisely the GRPO gradient we analyze in Sec. 2 and Eq. (9).
Difference to our surrogate reward.
A natural question arises: How does the PPO-style objective in Eq. (31) relate to the surrogate reward $2\arcsin(\sqrt{\rho_{\theta}})$ that Corollary 1 identifies as the function optimized by the GRPO gradient (Eq. (32))?
The resolution lies in recognizing that these are fundamentally different types of objectives:
  • As mentioned, the PPO-style objective (Eq. (31)) is defined over a fixed distribution $\pi_{\theta_{\mathrm{old}}}$.
  • Our surrogate reward identifies the on-policy population objective that the gradient in Eq. (32) ascends.
To illustrate this distinction, consider what happens if we incorrectly interpret the PPO objective as fully on-policy by substituting $\pi_{\theta_{\mathrm{old}}}=\pi_{\theta}$ everywhere (both in the integrand and the sampling distribution) in Eq. (31). This would yield:
$\mathbb{E}_{y_{i}\sim\pi_{\theta}(\cdot|x)}\Big[\frac{1}{N}\sum_{i=1}^{N}A_{i}\Big]\,.$ (33)
However, recall that $A_{i}={(r_{i}-\hat{\rho})}/{\sqrt{\hat{\rho}(1-\hat{\rho})}}$: by construction, the advantages are zero-centered: $\sum_{i\in[N]}A_{i}=0$ for any batch. Therefore, this naive on-policy interpretation yields a trivial objective that is identically zero; clearly not what GRPO optimizes.
As established in Corollary 1, the actual population objective that the on-policy GRPO gradient (Eq. (32)) ascends is the surrogate reward: $2\arcsin(\sqrt{\rho_{\theta}})$.
Perspective Beyond the Fully Online Setting.
This surrogate reward perspective provides a principled way to interpret and design advantage functions. Our analysis shows how algorithms like GRPO and $\widetilde{\text{GRPO}}_{K}$ correspond to specific surrogate rewards in the on-policy setting. This link is constructive: new advantage-shaping methods, derived from new surrogate rewards (as in Sec. 3 and 6), are not limited to the fully online setting. Their resulting advantage scores, $A_{i}$, can be seamlessly embedded within the PPO-style objective (Eq. (30)), allowing them to be used in practical, multi-epoch training regimes.
Appendix B Missing Proofs
B.1 Auxiliary Technical Results